# The Galactic population of canonical pulsars II

### v1

@@ 3.5 Gamma-ray detection
Figure 21: $P-\dot{P}$ diagram of the gamma-only pulsars for both the simulations and the observations, with 1 million pulsars simulated.
@@
What if we run another simulation (with the same parameters as in Table 5), by only changing the threshold of detection for gamma-ray ? We did this in order to see how many more pulsars we would detect in gamma-ray and where they would be in the $P-\dot{P}$ diagram, and also to deal with the problematic of the GeV excess in the Galactic centre. Therefore we increased $F_{\rm min}$ by 10 : if the galactic latitudes of the pulsar is ¡ 2°, then $F_{\rm min}=0.4\times 10^{-15}$ W.m-2, and concerning blind searches $F_{\rm min}=1.6\times 10^{-15}$ W.m-2.
@@
Figure 22: Spatial distribution of the simulated gamma-ray detected pulsars, projected onto the Galactic plane for 1 million pulsars simulated, with a lower $F_{min}$.
@@
The gamma-ray detected pulsar positions in the X-Y plane of the Galaxy are shown in Fig. 22. In this simulation, 1832 gamma-ray pulsars are detected, among them only 168, in red in Fig. 22, would have been detected if the same sensitivity as before was kept. Therefore we detect 11 times more gamma-ray pulsars compared to the number found with the previous sensitivity. Furthermore, the other 1664 pulsars, in green in Fig. 22, detected with this increased sensitivity highlight several features. We obviously detect more pulsars close to us, and also many more in the centre of the Galaxy. This result indicates that the GeV excess in the Galactic centre could be linked to pulsars not yet identified, while usually the hypothesis is that the GeV excess corresponds to self-annihilating dark matter particles (Hooper, 2022). In addition, millisecond pulsars, not modeled in this work, could however also contribute to this GeV excess.
@@ 4 Discussion
For the initial spin period distribution Igoshev et al. (2022) suggested values for the mean period $P_{\text{mean}}$ in the range from approximately $57$ ms to $129$ ms and for the spread $\sigma_{p}$ the range from $0.45$ to $0.65$. The values $P_{\text{mean}}=129$ ms and $\sigma_{p}=0.45$ are best educated guesses. However it is obvious that further investigations could benefit from an automated method to find the best set of values. Hence, we are currently working on an optimization algorithm to find the optimal initial parameters for the population synthesis, not only for the spin period, but for the whole parameter space presented in Table 5, except for $\alpha_{d}$ and $k_{\tau_{d}}$. Our optimization algorithm will help to constrain the properties of the Milky Way pulsar population by constraining parameters like the birth rate which is supposed to be between $1/33$ yr-1 and $1/150$ yr-1, see Faucher-Giguère & Kaspi (2006), Gullon et al. (2014) or Johnston & Karastergiou (2017).
We also added the effect of the ISM on the radio pulse profile, rendering more difficult to observe pulsars farther away than $10$ kpc. Furthermore, because of this plasma propagation effect fewer radio pulsars are detected compared to the observations. If the simulations discard the ISM effect, much more pulsars are detected at distances larger than $10$ kpc and with higher $\dot{E}$ as well, improving the match between the simulation and the observation. Perhaps, the effect of the ISM is overestimated when computing the scattering and dispersion effect converted in equation (28a) with $\tau_{scat}$ and $\tau_{DM}$ is too important and should have less influence on the observed width pulse profile, therefore with formulas that would give lower $\tau_{DM}$ and $\tau_{scat}$, we might get more realistic results.
@@
In the main part of the paper, we abstain from showing outcomes stemming from simulations featuring a constant magnetic field. However we ran such simulations that clearly highlight the inferior performance of a simulation without magnetic field evolution as can be checked in appendix B. Moreover appendix C explores different parameter spaces for the PPS in order to meticulously scrutinizing whether the parameters showcased in the main part of the paper optimally encapsulate observational data. This demonstrates that even if a substantial number of parameters are required for the PPS, they can be meaningfully constrained by such studies with reasonable accuracy.
@@
Furthermore, in Subsection 3.5 we explore the results when the sensitivity in gamma-ray is increased by a factor of 10. However we did not show what would happen if we tried to make a prediction for the radio detection by increasing the sensitivity by a factor of 50 to match approximately the properties of SKA. A simulation was run with this increased sensitivity and 52276 pulsars were detected (when simulating 1 million pulsars, with a death line), therefore approximately 31 times more than with the sensitivity of Parkes and Arecibo telescopes.
@@ 5 Summary
Without the implementation of the Galactic potential, in the PPS of Dirson et al. (2022), the pulsars were moving in straight lines along one direction chosen randomly for each pulsar. This evolution of the proper motion was not realistic, even though the evolution model had already replicated a significant portion of the canonical pulsar population. However all the new improvements put into this PPS, meaning : the Galactic potential, the effect of the ISM on the radio pulse profiles and the fact that the decay timescale of the magnetic field is taken randomly within a certain range, allow a more accurate reproduction of the population in the $P-\dot{P}$ diagram.
@@
Nonetheless, simulating 1 million pulsars allowed to show that with or without death line the results were alike and therefore, in the case where the oldest pulsar simulated is $4.1\times 10^{7}$ yr, a death line is not really necessary to explain the observations. In this paper we showed two approach to reproduce the observed canonical pulsars population, however when we consider the approach where the oldest pulsar simulated is only $4.1\times 10^{7}$ yr we state that no older pulsars might exist in the Galaxy, while the Milky Way is very much older than this. There is apparently no reason to not simulate pulsars that are older than this, and if there is one we do not know it (they become harder to detect because of their wide pulse profile with their age, but some are still detectable). If this situation is the reality then it means two things : the age of the observed pulsars are overestimated (because we observe pulsars older than $4.1\times 10^{7}$ yr in the canonical population, even if they are not numerous) and the death line is not necessary in pulsar population synthesis work to explain what we observe. When the simulation is ran with 10 millions pulsars, the death line becomes a necessity. However this simulation gives results less similar to the observations, because of a few simulated data which do not match any observational data.
@@
It was also shown that the old pulsars with age larger than 10 Myr have a tendency to lose their alignment between rotation axis and velocity, a direct effect of the Galactic potential as shown by the observations of Noutsos et al. (2013), but it is the first time it is done in a simulation.
@@
The striped wind model seems to reproduce the gamma-ray population of canonical pulsars very well, however in the simulation a higher number of these pulsars are detected compared to reality. It probably indicates that a part of unidentified sources from the 4FGL catalog are pulsars. Testing an increased sensitivity for the detection of gamma-ray pulsars lead us to think that a part of the origin of the GeV excess in the Galactic centre could be young gamma-ray pulsars, and when the millisecond pulsars will be modeled in our simulation we would also be able to check that these pulsars might be responsible for this GeV excess. The usual picture is to suppose that only millisecond pulsars are responsible, while young gamma-ray pulsars could also play a role.
@@
In summary, the best parameters found for this pulsar population synthesis are for $P_{\text{mean}}=129$ ms, $\sigma_{p}=0.45$ for the spin period distribution at birth, $B_{\text{mean}}=2.75\times 10^{8}$ T, $\sigma_{b}=0.5$ for the magnetic field distribution at birth and for a birth rate of 1/41 yr-1. Ultimately, this work provides a simulated population of pulsars within the Milky Way more similar to the observations compared to Johnston et al. (2020) and Dirson et al. (2022), the two most recent population synthesis considering both radio and gamma-ray emission. Moreover, such investigations will be useful to make more predictions for future radio surveys such as SKA, or NenuFAR.
@@
The PPS method used in this study could be applied to model populations of pulsars in the Milky Way other than the canonical ones. The magnetars, for instance, would need : another distribution at birth for their magnetic field to be modeled correctly, taking into account the emission in X-rays and having a faster decay for the magnetic field. The magnetars are the pulsars just above the canonical pulsars in the $P-\dot{P}$ diagram, at the top right of the diagram. A more complex population is the recycled or millisecond pulsars, that were ”brought back to life” by accreting matter from a companion star. This accretion mechanism spins up the neutron stars moving them from the graveyard of pulsars to the bottom left of the $P-\dot{P}$ diagram, crossing again the death line in the opposite sense and allowing them to create electron-positron pair cascades again. However, they are more complex to model because this evolution scenario should take into account the accretion process and spin-up. We are also currently working on both populations to reproduce them in a PPS based on this work.
@@ A.1 The scheme
here ${r}$ is the position vector, ${v}$ is the velocity vector, $h$ is the time step, ${\nabla}\Phi$ is the gradient of the potential and $n$ is the time at which we compute the position or the velocity. $\lambda$, $\xi$ and $\chi$ are constants that are tabulated in Table 10, see for instance Peter Young (2014).
@@ Appendix B Results for a constant magnetic field
Hence, this simulation was run with 1 million pulsar simulated, without a death line and without a decaying of the magnetic field. The KS test gave p-values of $6.3\times 10^{-22}$ and $9.8\times 10^{-77}$, therefore the null hypothesis is rejected for both $P$ and $\dot{P}$. The parameters used are in Table 5. It is less good than with a decaying magnetic field. As can be seen in Fig. 25, too many pulsars from the simulation appear on the top right of the diagram and almost none at the bottom. In Table 11, we note a total of pulsars which is smaller from what we found for all of our previous results in Sect. 3.
@@ Appendix C Influence of the different parameters
In order to show that the parameters chosen in the main part of this paper are meticulously picked and that the uncertainties brought by the large number of parameters is manageable, several simulations were run with different values for the main parameters of the initial distributions, showing a deviation from the best values gives bad results. The simulations were run with in the same conditions as Subsection 3.2, with only one parameter changing between the different sets of simulation.
Table 12: Influence of the birth rate.
Birth rate (in yr-1) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
1/30 2388 1964 204 220
1/80 963 799 79 85
1/120 696 597 50 49
Table 13: Influence of the mean initial period.
$P_{\rm mean}$ (in ms) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
57 1499 1134 193 172
70 1623 1253 215 155
100 1652 1348 151 153
Table 14: Influence of the mean initial magnetic field.
$B_{mean}$ (in T) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
$10^{7}$ 8407 6943 406 1058
$5\times 10^{7}$ 4447 3646 306 495
$5\times 10^{8}$ $1$ 933 109 69
The $P-\dot{P}$ diagram of each simulation is not shown here because almost every simulation finds a too high (or too low) number of pulsars, therefore the $P-\dot{P}$ diagrams are showing large differences between observations and simulations. Nonetheless, the number of pulsars detected for two of the simulations with different $P_{\rm mean}$ (with $P_{\rm mean}=70$ ms and $P_{\rm mean}=100$ ms) gives number closer to reality. However their p-values for $P$ and $\dot{P}$ when using the KS test were significantly below $0.05$, which allowed us to reject the null hypothesis for the distributions of $P$ and $\dot{P}$, for the results shown in Table 12, 13 and 14.
Moreover in order to get an idea on a good range of values for the parameters, the d-value obtained thanks to the KS test are shown in Table 15, 16, 17 and 18. For each parameter, 10 simulations were made in order to get a mean and a standard deviation for the d-value associated. These simulations helped finding the best values for the parameters by considering the parameter is well constrained when its d-values for both distribution of $P$ and $\dot{P}$ are below, or really close to 0.05 (meaning 5 % similarity between the observations and simulation distribution). Even though it is not an optimisation (and therefore the results can be improved on that point), it allows us to conclude that $P_{\rm mean}$ should be between 120 ms and 135 ms (with 129 ms as best value). $B_{\rm mean}$ should be between $2.25\times 10^{8}$ T and $2.8\times 10^{8}$ T (with $2.75\times 10^{8}$ T as best value). The birth rate should be between 1/37 yr-1 and 1/45 yr-1 (with 1/41 yr-1 as best value). Finally, $\sigma_{b}$ is well constrained between 0.49 and 0.51 (with 0.5 as best value). Concerning $\sigma_{p}$, it always gives d-values between 0.04 and 0.057 when $\sigma_{p}$ is below 0.6, therefore the value 0.45 was chosen as suggested by Igoshev et al. (2022).
@@
$P_{\rm mean}$ (in ms) d-value ($\dot{P}$) d-value ($P$)
90 $0.072\pm 0.0073$ $0.0577\pm 0.0080$
100 $0.0651\pm 0.00692$ $0.0503\pm 0.00636$
110 $0.0652\pm 0.00688$ $0.0381\pm 0.0074$
120 $0.0554\pm 0.00714$ $0.0302\pm 0.00468$
129 $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
135 $0.0544\pm 0.00768$ $0.0329\pm 0.00767$
150 $0.0572\pm 0.00779$ $0.0388\pm 0.00935$
@@
$B_{\rm mean}$ (in T) d-value ($\dot{P}$) d-value ($P$)
 $1\times 10^{8}$ $0.113\pm 0.00823$ $0.203\pm 0.00675$
$2\times 10^{8}$ $0.0418\pm 0.00459$ $0.0658\pm 0.00705$
$2.25\times 10^{8}$ $0.0456\pm 0.00455$ $0.0455\pm 0.0061$
$2.75\times 10^{8}$ $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
$2.8\times 10^{8}$ $0.0576\pm 0.00907$ $0.0298\pm 0.00728$
$3\times 10^{8}$ $0.0648\pm 0.01092$ $0.0426\pm 0.00775$
$3.25\times 10^{8}$ $0.0761\pm 0.00930$ $0.0480\pm 0.00997$
@@
$\sigma_{b}$ d-value ($\dot{P}$) d-value ($P$)
0.45 $0.0767\pm 0.00668$ $0.0513\pm 0.01258$
0.48 $0.0621\pm 0.00880$ $0.0385\pm 0.01234$
0.49 $0.0646\pm 0.00974$ $0.0304\pm 0.00501$
0.5 $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
0.51 $0.056\pm 0.00735$ $0.0306\pm 0.00701$
0.52 $0.0547\pm 0.00786$ $0.0358\pm 0.00579$
@@
BR (in yr -1) d-value ($\dot{P}$) d-value ($P$)
1/33 $0.0611\pm 0.01025$ $0.0313\pm 0.00636$
1/37 $0.0565\pm 0.00851$ $0.0295\pm 0.00645$
1/41 $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
1/45 $0.0585\pm 0.00510$ $0.0302\pm 0.00565$
1/70 $0.0691\pm 0.00830$ $0.0342\pm 0.01165$
1/150 $0.1165\pm 0.01564$ $0.0408\pm 0.00671$

### v2

@@ 3.5 $\gamma$-ray detection
Figure 20: $P-\dot{P}$ diagram of the $\gamma$-only pulsars for both the simulations and the observations, with 1 million pulsars simulated.
Figure 21: Distribution of the $\gamma$-ray flux of the simulated population along with the observations from the 3PC catalogue.
@@
To conclude the part about $\gamma$-ray pulsars, we predict the number of pulsars that would be detected by a future instrument, ten times more sensitive than Fermi/LAT. We asked the question: what if we run another simulation with the same parameters as in Table 5, only changing the threshold of detection for $\gamma$-ray? This investigation could help in identifying the nature of the GeV excess in the Galactic centre. Therefore we decreased the instrument sensitivity $F_{\rm min}$ by a factor 10: if the galactic latitudes of the pulsar is ¡ 2°, then $F_{\rm min}=0.4\times 10^{-15}$ W.m-2, and concerning blind searches we set $F_{\rm min}=1.6\times 10^{-15}$ W.m-2.
@@
Figure 22: Spatial distribution of the simulated $\gamma$-ray detected pulsars, projected onto the Galactic plane for 1 million pulsars simulated, with a lower $F_{min}$.
@@
The $\gamma$-ray detected pulsar positions in the X-Y plane of the Galaxy are shown in Fig. 22. In this simulation, 2492 $\gamma$-ray pulsars are detected, in green, being $\gamma$-ray only or radio-loud $\gamma$-ray pulsars, among them only 350 (really close to the number of $\gamma$-ray and radio/$\gamma$-ray pulsars detected in Subsection 3.2) would have been detected if the same sensitivity as before was kept, in red in Fig. 22. Therefore we detect 7 times more $\gamma$-ray pulsars compared to the number found with the previous sensitivity. Furthermore, the other 2142 pulsars, in green in Fig. 22, detected with this increased sensitivity highlight several interesting features. We obviously detect more pulsars close to Earth, and also many more in the centre of the Galaxy. This result indicates that the GeV excess in the Galactic centre could be linked to pulsars not yet identified, while usually the origin of this GeV excess is attributed to self-annihilating dark matter particles (Hooper, 2022). However, millisecond pulsars, not modeled in this work, could also contribute significantly to this GeV excess.
@@ 4 Discussion
For the initial spin period distribution Igoshev et al. (2022) suggested values for the mean period $P_{\text{mean}}$ in the range from approximately $57$ ms to $129$ ms and for its spread $\sigma_{p}$ the range from $0.45$ to $0.65$. The values $P_{\text{mean}}=129$ ms and $\sigma_{p}=0.45$ are best educated guesses. However further investigations on this work could benefit from an automated method to find the best set of values. Hence, we are currently working on an optimization algorithm to find the optimal initial parameters for the population synthesis, not only for the spin period, but for the whole parameter space presented in Table 5. Our optimization algorithm will help to constrain the properties of the Milky Way pulsar population by constraining parameters like the birth rate which is supposed to be between $1/33$ yr-1 and $1/150$ yr-1, see Faucher-Giguère & Kaspi (2006), Gullon et al. (2014) or Johnston & Karastergiou (2017).
@@
In the main part of the paper, we abstain from showing outcomes stemming from simulations featuring a constant magnetic field. However we ran such simulations that clearly highlight the inferior performance of a simulation without magnetic field evolution as can be checked in appendix B. Moreover appendix C explores different parameter spaces for the PPS in order to meticulously scrutinizing whether the parameters showcased in the main part of the paper optimally encapsulate observational data. This work demonstrates that even if a substantial number of parameters are required for the PPS, they can be meaningfully constrained by such studies with reasonable accuracy.
@@
Furthermore, in Subsection 3.5 we explored the results when the sensitivity in $\gamma$-ray is increased by a factor of 10. However we did not discuss improvements in the radio detection yet, while we run a simulation by increasing the sensitivity by a factor of 50 to match approximately the properties of SKA (McMullin et al., 2020). The result of this simulation with an increased sensitivity led to 52276 pulsars detected (when simulating 1 million pulsars, with a death line), therefore approximately 31 times more than with the sensitivity of Parkes and Arecibo telescopes.
We emphasize that the death valley strongly depends on the chosen model. For instance the radio luminosity law plays a central role, we assumed that $F_{r}\propto\dot{E}^{\alpha}$ with $\alpha=1/4$ (see equation (27)) according to Johnston & Karastergiou (2017) and Johnston et al. (2020) where they found it to best fit to the data. Following their conclusion we decided not to choose $\alpha$ as another parameter. However it would impact the death valley, but this effect is not considered in our work. Furthermore, in other works like Graber et al. (2024) for instance, they do not explicitly use a death valley because they use a different radio luminosity law, meaning a different value for $\alpha$ and adjust the magnetic field decay time accordingly. We stress that this is just another method to deal with the bottom right corner of the $P-\dot{P}$ diagram, where pulsars are likely to stop emitting photons, since the death valley removes pulsars based on a combination of $P$ and $\dot{P}$. Graber et al. (2024) PPS is based on a phenomenological approach where the radio luminosity is tuned to conform to observations. In our approach, adding a death valley is based on a more physical ground related to the microphysics close to the neutron star polar cap.
@@ 5 Summary
Without the implementation of the Galactic potential, in the PPS of Dirson et al. (2022), the pulsars were moving in straight lines along one direction chosen randomly for each pulsar. This evolution of the proper motion was not realistic, even though the evolution model had already replicated a significant portion of the canonical pulsar population. However all the new improvements put into this PPS, meaning the Galactic potential, the effect of the ISM on the radio pulse profiles and the fact that the decay timescale of the magnetic field is taken randomly within a certain range, allow a more accurate reproduction of the population in the $P-\dot{P}$ diagram.
@@
Nonetheless, simulating 1 million pulsars allowed to show that with or without a death line the results were alike and therefore, in the case where the oldest simulated pulsar is $4.1\times 10^{7}$ yr (its real age), a death line is not necessary to explain the observations. In this paper we showed two approaches to reproduce the observed canonical pulsar population, however when we consider the approach where the oldest pulsar simulated is only $4.1\times 10^{7}$ yr we state that no older pulsars might exist in the Galaxy, while the Milky Way is very much older than this. There is apparently no reason not to simulate pulsars that are older than this, and if there is one we do not know it, they become harder to detect because of their wide pulse profile with their age, but some are still detectable. If this situation is the reality then it means that the death line is not necessary in pulsar population synthesis to explain observations. When the simulation is ran with 10 millions pulsars, the death line becomes a necessity. However this simulation gives results less similar to observations.
@@
It was also shown that old pulsars with age larger than 10 Myr have a tendency to lose their alignment between rotation axis and velocity, a direct effect of the Galactic potential as shown by the observations of Noutsos et al. (2013), but it is the first time it is done in a simulation.
@@
The striped wind model seems to reproduce the $\gamma$-ray population of canonical pulsars well according to the p-value of the KS test greater than 0.05 for $\dot{P}$. However the KS test gives a p-value lower than 0.05 for $P$, but there are several observed radio-loud $\gamma$-ray pulsars that the simulation struggles to reproduce (not only the ones with high $P$) and might explain this low p-value. However in the simulation a higher number of these pulsars are detected compared to observations. It probably indicates that a part of the unidentified sources from the 4FGL catalogue are pulsars. Testing an instruments with improved sensitivity for the detection of $\gamma$-ray pulsars lead us to the conclusion that part of the GeV excess in the Galactic centre could be young $\gamma$-ray pulsars. Moreover the assumption that millisecond pulsars could also contribute to this excess will be checked once a PPS model for millisecond pulsars will be available. The most common picture suppose that only millisecond pulsars are responsible for this excess, while we demonstrated that young $\gamma$-ray pulsars could also play a significant role.
@@
In summary, the best parameters found for this pulsar population synthesis are for $P_{\text{mean}}=129$ ms, $\sigma_{p}=0.45$ for the log-normal spin period distribution at birth, $B_{\text{mean}}=2.75\times 10^{8}$ T, $\sigma_{b}=0.5$ for the magnetic field distribution at birth and for a birth rate of 1/41 yr-1. Ultimately, this work provides a simulated population of pulsars within the Milky Way more similar to the observations compared to Johnston et al. (2020) and Dirson et al. (2022), the two most recent population synthesis considering both radio and $\gamma$-ray emission. Moreover, such investigations are useful to predict the detection rate of future radio surveys such as SKA or NenuFAR.
@@
The PPS method developed in this paper can be extended to other pulsar populations in the Milky Way such as millisecond pulsars and magnetars. The magnetars, for instance, would need another distribution at birth for their magnetic field, taking into account the emission in X-rays and having a faster decay for the magnetic field. The magnetars are located just above the canonical pulsars in the $P-\dot{P}$ diagram, at the top right of the diagram. A more complex population is the recycled or millisecond pulsars, that were ”brought back to life” by accreting matter from a companion star. This accretion mechanism spins up the neutron stars moving them from the graveyard of pulsars to the bottom left of the $P-\dot{P}$ diagram, crossing again the death line in the opposite sense and allowing them to revive by creating electron-positron pair cascades again. However, they are difficult to model because their evolution scenario must take into account their accretion phase and spin-up. We currently work on both these populations to reproduce them in a PPS based on this work.
@@ A.1 The scheme
here ${r}$ is the position vector, ${v}$ is the velocity vector, $h$ is the time step, ${\nabla}\Phi$ is the gradient of the potential and $n$ is the time at which we compute the position or the velocity. $\lambda$, $\xi$ and $\chi$ are constants that are tabulated in Table 10, see for instance Young (2014).
@@ Appendix B Results for a constant magnetic field
Hence, this simulation was run with 1 million pulsar simulated, without a death line and without a decaying of the magnetic field. The KS test gave p-values of $1.1\times 10^{-14}$ and $1.5\times 10^{-76}$, therefore the null hypothesis is rejected for both $P$ and $\dot{P}$. The parameters used are in Table 5. It gives results less close to the observations than with a decaying magnetic field. As can be seen in Fig. 25, too many pulsars from the simulation appear on the top right of the diagram and almost none at the bottom. In Table 11, we note a total of pulsars which is smaller from what we found for all of our previous results in Sect. 3. We could expect to reproduce equally well the population of observed pulsars with a constant magnetic field, however compared to the study of Johnston & Karastergiou (2017) where they used this assumption for the magnetic field, we do not evolve the braking index. In our simulations, we have $\dot{\alpha}<0$, $\dot{B}<0$ and $\dot{n}=0$ whereas in their simulation $\dot{\alpha}<0$, $\dot{B}=0$ and $\dot{n}\neq 0$. Having either the braking index or the magnetic field evolve, allows to reproduce well the population of pulsars.
@@ Appendix C Influence of the different parameters
This appendix shows that the parameters chosen in the main part of this paper are meticulously picked and that the uncertainties brought by the large number of parameters is manageable. Multiple simulations were conducted by varying parameter values for the initial distributions, demonstrating that deviations from the optimal values lead to poor results. The simulations were run with in the same conditions as subsection 3.2, with only one parameter changing between the different sets of simulations and without implementing the ISM, because it would be too time consuming to perform all these simulations. As a consequence in the tables from 12 to 16), the p-values are low, while when including the ISM effect they become larger. Therefore, the results and conclusions presented in this section are based on this simplification, which was primarily used to guide the choice of parameters. While these assumptions have helped to identify efficient parameters, it is important to stress that the proposed values are not necessarily optimal. They serve primarily as a foundation for refining the parameters under more realistic conditions. That is why before validating a parameter we checked that the p-values were satisfactory when including the ISM. We found each time a higher p-values.
In order to get an idea on a good range of values for the parameters, the d-values and p-values obtained thanks to the KS test are shown in Table 12, 13, 14, 15 and 16. For each parameter, 10 simulations were performed in order to get a mean and a standard deviation for the d-values and p-values associated. These simulations helped finding the best values for the parameters by considering the parameter is well constrained when its d-values for both distribution of $P$ and $\dot{P}$ are below or close to 0.05 (meaning 5% similarity between the observed and simulated distributions), especially if the previous and next parameter values have higher d-values and lower p-values. As usual p-values above 0.05 means that the null hypothesis can not be rejected, however as explained above, without the ISM effect taken into account the p-values for the $\dot{P}$ distribution do not get too high here. Even though it is not an optimisation (and therefore the results can be improved on that point), it allows us to conclude that $P_{\rm mean}$ must be between 120 ms and 135 ms, with 129 ms as best value. $B_{\rm mean}$ must be between $2.25\times 10^{8}$ T and $2.8\times 10^{8}$ T, with $2.75\times 10^{8}$ T as best value. $\sigma_{b}$ is well constrained between 0.49 and 0.51, with 0.5 as best value. While $\sigma_{p}$ must be between 0.42 and 0.47, with 0.45 as best value. Finally, the birth rate must be between 1/37 yr-1 and 1/45 yr-1, with 1/41 yr-1 as best value.
@@
$P_{\rm mean}$ (in ms) d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
90 $0.072\pm 0.0073$ $0.058\pm 0.0080$ $0.0091\pm 0.021$ $0.038\pm 0.049$
100 $0.065\pm 0.0069$ $0.050\pm 0.0064$ $0.012\pm 0.0089$ $0.22\pm 0.0086$
110 $0.065\pm 0.0069$ $0.038\pm 0.0074$ $0.0033\pm 0.0022$ $0.32\pm 0.0076$
120 $0.055\pm 0.0071$ $0.030\pm 0.0047$ $0.0049\pm 0.0060$ $0.47\pm 0.25$
129 $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
135 $0.054\pm 0.0077$ $0.033\pm 0.0077$ $0.012\pm 0.015$ $0.091\pm 0.0095$
150 $0.057\pm 0.0078$ $0.039\pm 0.0093$ $0.011\pm 0.024$ $0.013\pm 0.0074$
@@
$B_{\rm mean}$ (in T) d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
 $1\times 10^{8}$ $0.11\pm 0.0082$ $0.20\pm 0.0067$ $4.4\times 10^{-16}\pm 1.2\times 10^{-15}$ $8.9\times 10^{-40}\pm 2.4\times 10^{-39}$
$2\times 10^{8}$ $0.042\pm 0.0046$ $0.066\pm 0.0070$ $0.033\pm 0.043$ $0.0058\pm 0.0072$
$2.25\times 10^{8}$ $0.046\pm 0.0045$ $0.045\pm 0.0061$ $0.011\pm 0.019$ $0.12\pm 0.076$
$2.75\times 10^{8}$ $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
$2.8\times 10^{8}$ $0.058\pm 0.0091$ $0.030\pm 0.0073$ $0.0074\pm 0.0055$ $0.20\pm 0.16$
$3\times 10^{8}$ $0.065\pm 0.011$ $0.043\pm 0.0077$ $8.7\times 10^{-4}\pm 9.0\times 10^{-4}$ $0.030\pm 0.050$
$3.25\times 10^{8}$ $0.076\pm 0.0093$ $0.048\pm 0.0010$ $0.0040\pm 0.0074$ $0.0019\pm 0.0034$
@@
$\sigma_{b}$ d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
0.45 $0.077\pm 0.0067$ $0.051\pm 0.013$ $7.2\times 10^{-4}\pm 0.0011$ $0.0042\pm 0.0078$
0.48 $0.062\pm 0.0088$ $0.038\pm 0.0123$ $0.0063\pm 0.0071$ $0.0063\pm 0.0071$
0.49 $0.065\pm 0.0097$ $0.030\pm 0.0050$ $0.059\pm 0.057$ $0.16\pm 0.02437$
0.5 $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
0.51 $0.056\pm 0.0073$ $0.031\pm 0.0070$ $0.0076\pm 0.0067$ $0.30\pm 0.18$
Table 15: D-value for different $\sigma_{p}$.
$\sigma_{p}$ d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
0.4 $0.055\pm 0.0024$ $0.024\pm 0.0056$ $0.0036\pm 0.0050$ $0.57\pm 0.24$
0.42 $0.054\pm 0.0055$ $0.028\pm 0.0051$ $0.0050\pm 0.0048$ $0.38\pm 0.23$
0.45 $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
0.47 $0.053\pm 0.0049$ $0.034\pm 0.0092$ $0.0096\pm 0.015$ $0.26\pm 0.21$
0.48 $0.053\pm 0.0045$ $0.037\pm 0.0084$ $0.0060\pm 0.0050$ $0.17\pm 0.18$
0.52 $0.053\pm 0.0049$ $0.043\pm 0.011$ $0.0082\pm 0.012$ $0.11\pm 0.14$
@@
BR (in yr -1) d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
1/33 $0.061\pm 0.010$ $0.031\pm 0.0064$ $0.013\pm 0.014$ $0.49\pm 0.17$
1/37 $0.056\pm 0.0085$ $0.029\pm 0.0064$ $0.0086\pm 0.0095$ $0.48\pm 0.21$
1/41 $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
1/45 $0.058\pm 0.0051$ $0.030\pm 0.0056$ $0.0055\pm 0.0042$ $0.18\pm 0.14$
1/70 $0.069\pm 0.0083$ $0.034\pm 0.0116$ $4.3\times 10^{-4}\pm 0.0012$ $0.048\pm 0.074$
1/150 $0.12\pm 0.016$ $0.041\pm 0.0067$ $5.9\times 10^{-12}\pm 1.2\times 10^{-11}$ $0.035\pm 0.091$